rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#2025 — FS#6000 — network switch 46.105.110.0/24 et 111.0/24

Attached to Project— Dedicated servers
Maintenance
RBX4
CLOSED
100%
Following the task http://status.ovh.net/?do=details&id=2020, we will
back to normal status tomorrow morning between 5 and 7 o'clock.

Several but very short network disconnection will take place.

Date:  Friday, 28 October 2011, 12:36PM
Reason for closing:  Done
Comment by OVH - Friday, 28 October 2011, 10:35AM

The works have been started.


Comment by OVH - Friday, 28 October 2011, 10:50AM

The simple recommended reboot by cisco did not restore the functioning.
Once again We have cut the .248 but the network does not respond to ping at all.

Currently We will replace the two devices.


Comment by OVH - Friday, 28 October 2011, 10:53AM

The replacement has been very long because of physical complications but it's over.
The conf has been implemented in the two switches, we check that everything is in redundant mode


Comment by OVH - Friday, 28 October 2011, 11:14AM

We have had problems with the SLA server which did not respond to ping
on the network so it announced that the servers are
down while they have been responding to ping for several minutes already.

This problem is fixed and all servers are UP.

We check again.


Comment by OVH - Friday, 28 October 2011, 11:16AM

The servers with ASA firewall have been still impacted by the switch.

Fixed.


Comment by OVH - Friday, 28 October 2011, 11:21AM

The original problem is still present, we look once again on how to stabilize the infrastructure.


Comment by OVH - Friday, 28 October 2011, 11:23AM

We are still on the problem.


Comment by OVH - Friday, 28 October 2011, 11:45AM

Trying to simplify the configuration in many ways,we did not succeed.
we have restarted everything from 0 starting from the backup and after a reboot of the switches,
it works.

The problem is back few minutes after returning to normal status, we are still in doubt.


Comment by OVH - Friday, 28 October 2011, 12:22PM

Last step:restore the redundancy links to both routers.
We have just re-enable the 2nd link to the main router.


Comment by OVH - Friday, 28 October 2011, 12:36PM

The restoration of this link has been well achieved.

The balance is that the problem came from the VPC. A reboot of both devices at the
same time was needed.
The only way was to reset the original configuration and reboot the two switches at the same time.

We will probably do so next time: Following a loss of one of the
VPs equipment ,We MUST reboot the two devices to find
redundancy without problems. We will talk about this problem to cisco